OpenStudy (anonymous):

write an equation and solve for the variable specified. Five more than a number g is six less than twice a number h. Solve for g.

OpenStudy (anonymous):

(x-c)/2 = d x-c = 2 x = c+2d ::::: " Five more than a number g is six less than twice a number h." 5 + g = 2h - 6 g = 2h - 11 ::::: "One fourth of a number q is three more than three times a number w." q/4 = 3 + 3w q = 12 + 12w ::::: "Eight less than a number s is three more than four times a number t." s - 8 = 3 + 4t s = 11 + 4t

OpenStudy (anonymous):

thanks joker2.0 you are the best

OpenStudy (anonymous):

welcome